A.   Determine how much vinegar in gal. is required to react with an iron nail to produce 32.5 g of hydrogen. The density of HC2H3O2 is 1.053 g/ml.


1
Expert's answer
2021-11-24T12:21:22-0500

2Fe + 6HC2H3O2 → 2Fe(C2H3O2)3 + 3H2


Moles of hydrogen = "\frac{32.5}{2}=16.25moles"


Moles of HC2H3O2 "=\frac{6}{3}×16.25=32.5moles"


Mass "=60\u00d732.5=1950grams"


Volume "=\frac{1950}{1.053}=1851.9ml"